# GOVERN: Gradient Orientation Vote Ensemble for Multi-Teacher Reinforced Distillation

### v1

@@ Abstract
Pre-trained language models have become an integral component of question-answering systems, achieving remarkable performance. For practical deployment, it is critical to carry out knowledge distillation to preserve high performance under computational constraints. In this paper, we address a key question: given the importance of unsupervised distillation for student performance, how does one effectively ensemble knowledge from multiple teachers at this stage without the guidance of ground-truth labels? We propose a novel algorithm, GOVERN, to tackle this issue. GOVERN has demonstrated significant improvements in both offline and online experiments. The proposed algorithm has been successfully deployed in a real-world commercial question-answering system.
@@ 1 Introduction
Traditional search engine aims at deliver relevant web pages to satisfy users’ question, while sometimes the single paragraph that answer the question might buried deep in a web page, it asks for a web-based Open domain Question Answering (OpenQA) system to find that needle-in-a-haystack information (e.g. Qu et al., 2021; Zhang et al., 2023).
BERT-liked pre-trained language models have achieved state-of-the-art performance in OpenQA (e.g. Zhang et al., 2021). However, due to computational costs, the direct application of these models in real-time search engines like Google is currently unfeasible. For instance, the top-performing models on the Natural Question dataset, R2-D2 (Fajcik et al., 2021) and UnitedQA (Cheng et al., 2021) come with 1.29 billion and 2.09 billion model parameters, respectively. Further complicating matters is the fact that ensemble methods, which can enhance performance, entail even greater computational overheads.
The distillation of knowledge from multiple teachers has emerged as a powerful technique for improving the performance and generalization of deep neural networks while reducing the computational cost. This two-stage training paradigm, which training large model with limited labeled data as teacher and then using it to generate soft label on large amount unlabeled data for the purpose of student training, was first proposed by Hinton et al. (2015). Since the knowledge from single teacher may be biased and inaccurate, ensemble distillation from multiple teachers was considered by previous works to achieve more promising performance (e.g. You et al., 2017; Fukuda et al., 2017a).
@@
In this paper, we propose an algorithm called Gradient Orientation Vote Ensemble Reinforced distillatioN (GOVERN) to do sample-wise dynamic teacher selection without the need of label guidance. In the stage of unsupervised distillation, we take the gradient orientation provided by teachers as vote and follow the majority rule to do sample-wise dynamic teacher selection. While in supervised distillation stage, we use GOVERN to improve confidence aware ensemble method proposed by Zhang et al. (2022) with the help of human label.
@@
  • We propose GOVERN to do sample-wise dynamic teacher selection without the need of label guidance. We also give a mathematical proof that GOVERN can perform better than mean ensemble. To the best of our knowledge, GOVERN is the first method which can find sample-wise high-quality teachers without label guidance.
  • We propose a distillation framework for industrial application which combines GOVERN into both of the unsupervised and supervised distillation stage.
  • We conduct extensive experiments to demonstrate the effectiveness of GOVERN. Results show that GOVERN is benefit in both distillation stage and can boost a real-world question answering system’s performance.
@@ 3.1 Unsupervised Distillation
While these weighting methods do account for the performance differences among various teacher models, they employ a uniform weighting coefficient for all samples during the distillation process. This approach neglects the varying emphasis on each teacher’s abilities and their respective confidence levels regarding samples with differing distributions.
@@ 3.1.1 GOVERN
The voted result is calculated as the sum of $Grad_{i}$:
@@
3.1.2 Proof
In this section, we mathematically prove that the sample-wise dynamic weighting ensemble algorithm GOVERN is better than mean-ensemble.
First, let’s consider the positive samples. For a binary classification model with an accuracy rate of p, the probability of correct classification after each sampling follows a Bernoulli distribution. Thus, the expected classification accuracy of a single teacher is p, and the variance is p(1-p).
To simplify computation, we assume the performance of the N teachers is consistent, i.e., $p=p_{1}=...=p_{N}$, where $p_{i}$ is the accuracy of $T_{i}$. The mean ensemble of N teachers is formulated as:
$X_{ME}=\frac{\Sigma_{i}X_{i}}{N}$ (16)
Therefore, compared to a single model, the mean-ensemble demonstrates better robustness.
$p_{0}=\sum_{m=\frac{N+1}{2}}^{N}C_{N}^{m}p^{m}(1-p)^{N-m}$ (17)
Utilizing mathematical induction, it is trivial to prove when $p>1/2$, $p_{0}>p$. This is called Condorcet’s jury theorem and details of proof can be found in (Sancho, 2022).
In the stage of distillation, with the assumption that teacher’s performance is better than student, we have $P(logit_{i}>logit_{0})>1/2$ for positive samples, where $logit_{i}$ is the output of $T_{i}$ and $logit_{0}$ is the output of student. So GOVERN is better than mean-ensemble.
For the negative samples, we can prove GOVERN is better in the same way.
@@ 3.2 Supervised Distillation: GOVERN-CA
According to the aforementioned formulas, we only select teacher with the correct gradient descent orientation. Besides, the teacher whose prediction closely align with the ground-truth labels is assigned a greater weight $\omega_{i}$. This weighting is attributed to the model’s substantial confidence in making accurate judgments, thereby providing correct guidance.
@@
Mean Ensemble Distillation on unlabeled data 125M 12 768 12 82.04% 74.63%
LR Ensemble Distillation on unlabeled data 125M 12 768 12 81.98% 75.24%
GOVERN on unlabeled data 125M 12 768 12 83.65% 76.02%
+ CA-MKD on labeled data 125M 12 768 12 82.68% 75.67%
+ GOVERN-CA on labeled data 125M 12 768 12 83.69% 76.43%
Table 2: Results of offline experiments. Metrics denoted in bold represent the best results in the unsupervised distillation phase, while underscored and bolded denote the best results in the supervised distillation phase.
@@ 4.2 Experiment Details
Teacher Architecture In order to obtain multiple models with different structure and ability, we use pretrained models with different layer and fine-tune them on different samplings of the total labeled data. The specific structural parameters for each teacher model can be found in Table 2. Each model has been trained using a sample of 90% of the total data for training purposes.
Student Architecture Considering the computing resources and time consuming, we use the 12-layer transformer structure as the model for online deployment.
@@ 4.3 Evaluation Metrics
Precision-Recall is a useful measure of success of prediction when the classes are very imbalanced. $P=T_{p}/(T_{p}+F_{p}),R=T_{p}/(T_{p}+F_{n})$ ,where $T_{p}$, $F_{p}$ and $F_{n}$ represent for the number of true positives, false positives and false negatives, respectively.
@@ 4.4 Main Results
Logistic Regression Weighted Ensemble We trained a logistic regression model based on the development set to determine the weighting coefficients, and use them to obtain a weighted ensemble model.
CA-MKD This an algorithm proposed by Du et al. (2020) which adaptively assigns sample-wise reliability for each teacher prediction with the help of ground-truth labels, with those teacher predictions close to one-hot labels assigned large weights.
@@ 4.6 Ablation Study
We executed a series of experiments to illustrate the applicability of our method under various conditions. Due to computational resource constraints, we employed a 12-layer transformer as the teacher model and a 4-layer transformer as the student model for ablation study. The training data was divided into ten folds, with each teacher model trained on nine out of the ten folds, resulting in ten distinct teacher models. For the distillation process, we utilized fifty million unlabeled samples. The epoch for training was set to one.
@@ 5 Related Work
Unsupervised Ensemble Distillation There are a few works focused on the ensemble method on unsupervised data Li and Wang, 2019; Sui et al., 2020, these works simply use the average prediction of multiple teachers as the distillation signal. Recently, Wu et al. (2022) and Iliopoulos et al. (2022) made efforts on knowledge distillation with unlabeled examples, but these studies primarily concentrate on dynamically assigning weight to unlabeled examples. These approaches do not address the issue of teachers specializing in varying sample distributions.
Supervised Ensemble Distillation Li et al. (2021) first explored the idea of dynamic knowledge distillation with the help of ground-truth label. Yuan et al. (2021) proposed a novel RL-based approach, which dynamically assigns weights to teacher models at instance level. Cai et al. (2022) proposed algorithm ensembles multi-teacher logits supervised by human-annotated labels in an iterative way. Zhang et al. (2022) introduced confidence-aware mechanism on both predictions and intermediate features for multi-teacher knowledge distillation.
@@ 6 Conclusion
In this paper, we present a novel algorithm, GOVERN, which dynamically selects teachers based on their gradient descent orientation. This algorithm does not require ground-truth labels, making it suitable for use during unsupervised distillation stages. Additionally, it can be integrated with existing supervised ensemble methods. The effectiveness of our method is affirmed through extensive experimentation.
@@
The proposed GOVERN algorithm does not currently account for the varying performance levels of different teachers. This could be a shortcoming as it may be beneficial to assign a higher weight to more competent teachers, even if they share the same gradient descent orientation as other selected teachers.

### v2

@@ Abstract
Pre-trained language models have become an integral component of question-answering systems, achieving remarkable performance. However, for practical deployment, it is crucial to perform knowledge distillation to maintain high performance while operating under computational constraints. In this paper, we address a key question: given the importance of unsupervised distillation for student model performance, how can knowledge from multiple teacher models be effectively ensemble during this stage without the guidance of labels? We propose a novel algorithm, GOVERN, to tackle this issue. GOVERN has demonstrated significant improvements in both offline and online experiments, enabling the student model to achieve results comparable to that of teacher ensembles. Our experiments show that GOVERN remarkably requires a mere 1% of the ensemble method’s inference budget to achieve 99.5% of performance. The proposed algorithm has been successfully deployed in a real-world commercial question-answering system, demonstrating its real-world applicability.
@@ 1 Introduction
Traditional search engine aims at deliver relevant web pages to satisfy users’ question, while sometimes the single paragraph that answer the question might buried deep in a web page, it asks for a web-based Open domain Question Answering (OpenQA) system to find that needle-in-a-haystack info (e.g. Qu et al., 2021; Zhang et al., 2023).
BERT-liked pre-trained language models have achieved state-of-the-art performance in OpenQA (e.g. Zhang et al., 2021). However, due to computational costs, the direct application of these models in real-time search engines like Google is currently unfeasible. For instance, the top-performing models on the Natural Question dataset, R2-D2 (Fajcik et al., 2021) and UnitedQA (Cheng et al., 2021) come with 1.29B and 2.09B model parameters. Further complicating matters is the fact that ensemble methods, which can enhance performance, entail even greater computational overheads.
The distillation of knowledge from multiple teachers has emerged as a powerful technique for improving the performance and generalization of DNN while reducing the computational cost. This two-stage training paradigm, which training large model with limited labeled data as teacher and then using it to generate soft label on large amount unlabeled data for the purpose of student training, was first proposed by Hinton et al. (2015). Since the knowledge from single teacher may be biased and inaccurate, ensemble distillation from multiple teachers was considered by previous works to achieve more promising performance (e.g. You et al., 2017; Fukuda et al., 2017a).
@@
In this paper, we propose Gradient Orientation Vote Ensemble Reinforced distillatioN (GOVERN) to do sample-wise dynamic teacher selection without the need of label guidance.
@@
  • We propose GOVERN to do sample-wise dynamic teacher selection without the need of label guidance. We also give a proof that GOVERN can perform better than mean ensemble. To the best of our knowledge, GOVERN is the first method which can find sample-wise high-quality teachers without label guidance.
  • We propose a novel distillation framework for industrial applications that integrates the GOVERN method into both unsupervised and supervised distillation stages. This framework enhances the performance of student neural networks, enabling them to achieve results comparable to those of ensemble methods. The potential benefits of this approach make it a valuable contribution to industrial OpenQA systems.
  • Extensive experiments show that GOVERN is benefit in both distillation stage and can boost the real-world question answering system.
@@ 3.1 Unsupervised Distillation
While these weighting methods do account for the performance differences among various teachers, they employ a uniform weighting coefficient for all samples during the distillation process. This approach neglects the varying emphasis on each teacher’s abilities and their respective confidence levels regarding different samples.
@@
It is noted that previous works like UniKD(Wu et al. (2022)) and wVID(Iliopoulos et al. (2022)) have explored the dynamic assignment of weights. But these methods are used to to evaluate the significance of unlabeled examples, rather than assessing the importance of teachers. These methods could be synergistically integrated with the GOVERN framework, as they enhance unsupervised distillation from distinct perspectives.
@@ 3.1.1 GOVERN
The voted result is calculated as:
@@
In Appendix A, we give a prove that the sample-wise dynamic weighting ensemble algorithm GOVERN is better than mean-ensemble.
@@ 3.2 Supervised Distillation: GOVERN-CA
Thereby, we only select teacher with the correct gradient descent orientation. Besides, the teacher whose prediction closely align with the ground-truth labels is assigned a greater weight $\omega_{i}$. This weighting is attributed to the model’s substantial confidence in making accurate judgments, thereby providing correct guidance.
@@
Mean Ensemble Distillation on unlabeled data 125M 12 768 12 82.04% (0.07) 74.63% (0.12)
LR Ensemble Distillation on unlabeled data 125M 12 768 12 81.98% (0.11) 75.24% (0.12)
GOVERN on unlabeled data 125M 12 768 12 83.65% (0.08) 76.02% (0.14)
+ CA-MKD on labeled data 125M 12 768 12 82.68% (0.03) 75.67% (0.05)
+ GOVERN-CA on labeled data 125M 12 768 12 83.69% (0.06) 76.43% (0.09)
Table 2: Results of offline experiments. Metrics denoted in bold represent the best results in the unsupervised distillation phase, while underscored and bolded denote the best results in the supervised distillation phase. All distilled results are average taken over 5 random seeds with standard deviation in parenthesis.
@@ 4.2 Experiment Details
Teacher Architecture In order to obtain multiple models with different structure and ability, we use the series of pretrained models ERNIE-2.0 (sun2020ernie) with different layer and fine-tune them on different samplings of the total labeled data. The specific structural parameters for each teacher model can be found in Table 2. Each model has been trained using a sample of 90% of the total data for training purposes.
Student Architecture Considering the computing resources and time consuming, we use the 12-layer transformer structure for online deployment.
@@ 4.3 Evaluation Metrics
Precision-Recall is a useful measure of success of prediction when the classes are very imbalanced. $P=T_{p}/(T_{p}+F_{p}),R=T_{p}/(T_{p}+F_{n})$ ,where $T_{p}$, $F_{p}$ and $F_{n}$ represent for the number of true positives, false positives and false negatives.
@@
r(query_change) The query change ratio, defined as the proportion of sessions where users initiate a subsequent search following their initial query, serves as an online user behavior metric. This study reports only the difference in the query change ratio between the experimental and baseline methods, withholding absolute values.
Lower query change ratio reflects better performance as users are satisfy with the initial response, obviating the necessity for further queries.
r(skip_click) The skip click ratio, quantified as the proportion of instances where users click on web pages below the answer card (figure 2), indicates potential dissatisfaction with the answer provided. Due to confidentiality constraints, we report only the differential in skip click ratios between the experimental and baseline methods.
Figure 2: The Answer Card is retrieve by the question answering system. Web pages below are not display in answer card format.
@@ 4.4 Main Results
Logistic Regression Weighted Ensemble We trained a logistic regression model based on a dev-set to determine the weighting coefficients, and use these to obtain the weighted-sum of scores.
MED(Mean Ensemble Distillation) The predict score produced by Mean Ensemble Teachers is used as the optimizing object of student.
LRED(LR Ensemble Distillation) This variant uses Logistic Regression Weighted Ensemble Teachers for distillation instead of Mean Ensemble.
CA-MKD This an algorithm proposed by Zhang et al. (2022) which adaptive assigns sample-wise reliability for each teacher prediction with the help of ground-truth labels, with those teacher predictions close to one-hot labels assigned large weights.
It is noted that the 125M distilled model outperforms the 10B teacher model. This could be attributed to the limited size of the training set, which comprises only 19K distinct questions and 2.5M labeled question-pair examples. Such a dataset may not be sufficiently large to leverage the full potential of the larger model. Additionally, an increase in the performance of the teacher models was noted throughout a finetuning epoch, suggesting that these models are underfitted.
@@ 4.5 Online Experiment
$\Delta_{query\_change}$0.68% -1.03%
$\Delta_{skip\_click}$3.46% -4.76%
@@ 4.6 Ablation Study
Due to computational resource limitations, our ablation study utilized a 12-layer transformer as the teacher model and a 4-layer transformer as the student model. We divided the training data into ten folds, training each of the ten distinct teacher models on nine folds. The distillation process involved fifty million unlabeled samples, with the training epoch set to one.
The metric we report in this section is qp prAUC. This metric computes the area under the precision-recall curve where precision-recall is computed based on every qp-pair. It gives an overall measurement of classification ability.
@@ 5 Related Work
Unsupervised Ensemble Distillation There are a few works focused on the ensemble method on unsupervised data (Li and Wang, 2019; Sui et al., 2020), these works simply use the average output of multiple teachers as the distillation signal. Recently, Wu et al. (2022) and Iliopoulos et al. (2022) made efforts on distillation with unlabeled examples, but these studies primarily concentrate on dynamically assigning weight to unlabeled data. These approaches do not address the issue of teachers specializing in varying sample distributions.
Supervised Ensemble Distillation The idea of dynamic knowledge distillation with the help of ground-truth label was first explored by Du et al. (2020) and Li et al. (2021). Yuan et al. (2021) proposed a novel RL-based approach, which dynamically assigns weights to teacher models at instance level. Cai et al. (2022) proposed algorithm ensembles multi-teacher logits supervised by human-annotated labels in an iterative way. Zhang et al. (2022) introduced confidence-aware mechanism on both predictions and intermediate features for multi-teacher knowledge distillation.
@@ 6 Conclusion
In this paper, we present a novel algorithm, GOVERN, which dynamically selects teachers based on their gradient descent orientation. It does not require ground-truth labels, making it suitable for unsupervised distillation stages. Additionally, it can be integrated with existing supervised ensemble methods. The effectiveness of our method is affirmed through extensive experimentation.
@@
The GOVERN algorithm does not currently account for the varying performance levels of teachers. This could be a shortcoming as it may be beneficial to assign a higher weight to more competent teachers, even if they share the same gradient descent orientation as other selected teachers.
As mentioned in section 3.1, existing dynamic methods are typically used to assign significance to samples, allowing GOVERN to integrate with them. We leave such integration as future work.
Theoretically, GOVERN is a general method that can be applied to other classification tasks. We conducted experiments specifically on the QA task because our team is responsible for the question-answering function in a search engine. We encourage readers to explore its application in different use cases.
Appendix A Appendix
In this section, we mathematically prove that the sample-wise dynamic weighting ensemble algorithm GOVERN is better than mean-ensemble. We only make the proof on positive samples, as for the negative samples, the proof process is the same due to the symmetry.
A.1 Discrete Situation
First, we consider the discrete case where each $teacher_{i}$ can be viewed as a classifier. For a binary classification model with precision of p, the probability of correct classification after each sampling follows a Bernoulli distribution. Thus, the expected classification precision of a single teacher is p, and the variance is p(1-p).
To simplify computation, we assume the performance of the N teachers is consistent, i.e., $p=p_{1}=...=p_{N}$, where $p_{i}$ is the precision of $T_{i}$. The mean ensemble of N teachers is formulated as:
$X_{ME}=\frac{\Sigma_{i}X_{i}}{N}$ (25)
Due to the fact $E(X_{ME})=E(X_{M_{i}})$ and $D(X_{ME})<D(X_{M_{i}})$, we conclude the following lemma:
Lemma 1. Compared to the prediction from single model, although the mean ensemble result demonstrates better robustness, it keeping the expected precision the same.
$p_{0}=\sum_{m=\frac{N+1}{2}}^{N}C_{N}^{m}p^{m}(1-p)^{N-m}$ (26)
Utilizing mathematical induction, it is trivial to prove when $p>1/2$, $p_{0}>p$. This is called Condorcet’s jury theorem and details of proof can be found in (Sancho, 2022). Now we can state the following lemma:
Lemma 2. In discrete situation, vote-ensemble shows higher expected precision compared with mean-ensemble.
A.2 Consecutive Situation
It is noted that in the setting of distillation, we take model as scorer rather than a simple classifier, and the output of the scorer is a float in $[0,1]$. The distribution of the output is subject to Beta distribution, which is the conjugate distribution of Bernoulli distribution. This assumption can also be empirically verified as Figure 4 shows.
Figure 4: Left part shows the distribution of our model’s output on test set, and right part shows the distribution of $Beta(19.0,3.0)$. We can see that the model’s output keep similar distribution with Beta function.
To simplify computation, we assume all teachers is subject to the same distribution, i.e., $X_{i}\sim B(b_{1},b_{2}),\forall i\in\{1,...,N\}$. Then we have:
$E(X_{ME})$ $=\frac{\sum E(X_{i})}{n}=\frac{b_{1}}{b_{1}+b_{2}}$ (27)
$D(X_{ME})$ $=\frac{\sum D(X_{i})}{n^{2}}$ (28)
$=\frac{(b_{1}*b_{2})}{n*(b_{1}+b_{2})^{2}*(b_{1}+b_{2}+1)}$ (29)
So Lemma 1 still holds in consecutive situation.
Next, we consider the case where N teachers calculate the ensemble scores by utilizing GOVERN method. We conduct numerical simulation using Monte-Carlo sampling to verify the superiority of GOVERN.
Figure 5:
We set 10 teachers with same distribution as $X_{i}\sim B(20.0,2.0),\forall i\in\{1,...,N\}$, and set student as $X_{0}\sim B(19.0,3.0)$. The number of simulation is set to 1M.
The simulation result is shown in figure 5. We can see that the expectation of mean-ensemble is same with single teacher’s output, while the variance is lower. This result is consist with Lemma 1. Under the setting of GOVERN, it shows higher expectation compare with mean-ensemble, and keeps comparable variance. This verifies that GOVERN can obtain a better score with high expectation for distillation, and keep comparable robustness like mean-ensemble.